OpenStudy (anonymous):

Evaluate sin(arctan0)?

OpenStudy (anonymous):

\[\sin (\tan^{-1} 0)\]

OpenStudy (anonymous):

let arc tan 0=x, \[\tan x=0=\tan 0,x=0,\sin x=\sin 0=?\]

OpenStudy (anonymous):

So would it just be 0?